Find the values of the trigonometric functions of t from the given information.
cos t = − 12/13
Terminal point of t is in Quadrant III
sin t =
tan t =
csc t =
sec t =
cot t =
Please be descirptive on how this is solved. Thank you

Answer by stanbon(75887)

Find the values of the trigonometric functions of t from the given information.
cos t = − 12/13
Terminal point of t is in Quadrant III where x is negative and y is negative
-----
By definition, cos(t) = x/r
So x = -12 and r = 13
-----------------------------
Solve for "Y"::
y = -sqrt[13^2 - 12^2] = -sqrt(25) = -5
-----
sin = y/r = -5/13
cos = x/r = -12/13
tan = y/x = -5/-12 = 5/12
-------------
csc = r/y = -13/5
sec = r/x = -13/12
cot = x/y = 12/5
---------------------
Cheers,
Stan H.
----------------------